what is the mass (weight) gain of a patient who received 1.8L in intravenous glucose solution IV with a density of 1.23g/mL ?
1
Expert's answer
2012-01-10T10:48:11-0500
The mass (weight) gain of a patient is equivalent mass of glucose solution (m).
m = pV ( p is density)
V = 1,8 L = 1800ml.
m = 1.23*1800 = 2215 g (or 2.214 kg)